# Direct Debiased Machine Learning via Bregman Divergence Minimization

### v1

@@ 1.1 Contents of This Study
  • Generalized Riesz regression. We formulate the Riesz representer estimation problem as a Bregman divergence minimization problem. Bregman divergence depends on a convex function specified by us and yields various estimation methods by changing its definition. For example, Riesz regression and covariate balancing propensity scores are special cases. We refer to this estimation procedure as generalized Riesz regression.
@@ 6.2 Maximum Likelihood Estimation
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BR}}_{g^{\mathrm{KL}}}\big(\alpha\big),$
@@
$\widehat{\text{BR}}_{g^{\mathrm{KL}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\operatorname{sign}(\alpha(X_{i}))\log|\alpha(X_{i})|-\log\alpha((1,Z_{i}))-\log\alpha((0,Z_{i}))\right).$
This corresponds to unnormalized KL (UKL) divergence minimization in density-ratio estimation (Sugiyama et al., 2011b), which generalizes Kullback-Leibler Importance Estimation. Procedure (KLIEP).
@@ 6.3 Empirical Balancing
$g^{\mathrm{E}}(\alpha)=|\alpha-1|\log\left|\alpha-1\right|-|\alpha|.$
@@
$\partial g^{\mathrm{E}}(\alpha)=\begin{cases}-\log\left(\alpha-1\right)&\text{if}\ \alpha<0\\ \log\left(\alpha-1\right)&\text{if}\ \alpha>1\end{cases}.$
@@
$\text{BR}_{g^{\mathrm{E}}}\big(\alpha\big)\coloneqq\mathbb{E}\Big[\Big(\log|\alpha(X)-1|+|\alpha(X)|\Big)-\log|\alpha((1,Z))-1|-\log|\alpha((0,Z))-1|\Big].$
@@
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BR}}_{g^{\mathrm{E}}}\big(\alpha\big),$
@@
$\widehat{\text{BR}}_{g^{\mathrm{E}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\Big(\log|\alpha(X_{i})-1|+|\alpha(X_{i})|\Big)-\log|\alpha((1,Z_{i}))-1|-\log|\alpha((0,Z_{i}))-1|\right).$
@@
As in Section 6.2, model the Riesz representer via the inverse propensity score. Let $r((d,Z))$ model the inverse propensity score, that is, $r((1,Z))$ models $r_{0}((1,Z))\coloneqq 1/\pi_{0}(Z)$ and $r((0,Z))$ models $r_{0}((0,Z))\coloneqq 1/(1-\pi_{0}(Z))$. Using this model, define
@@
Then
@@
If $r((1,Z))=1/\pi(X)$ and $r((0,Z))=1/(1-\pi(X))$, then we have $r((1,Z))-1=1/\big(r((0,Z))-1\big)$. Therefore, it holds that
@@ 7.2 Generalized Riesz Regression
With $g^{\text{KL}}(\alpha)=|\alpha-1|\log|\alpha-1|-|\alpha|$, the dual problem enforces moment balance between treated units and reweighted controls (automatic covariate balancing), aligning with entropy–balancing–style weights but learned through the Riesz objective.
@@ 9.2 Generalized Riesz Regression via Bregman Matching
which yields the classical LSIF used in Kanamori et al. (2009). While Chernozhukov et al. (2025) proposes Riesz regression for covariate shift adaptation, the optimization problem is the same as LSIF, and the overall covariate shift adaptation method coincides with Kanamori et al. (2009). Kato et al. (2024) also proposes a doubly robust form for covariate shift adaptation by extending Kanamori et al. (2009).
@@ 10.1 Linear Models
where $\Phi\colon\{1,0\}\times{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a set of basis functions. For this model, using the squared loss (Riesz regression) automatically attains covariate balancing, as discussed in Bruns-Smith et al. (2025).
@@
where $\bm{0}_{n}$ is the $p$-dimensional zero vector. This optimization problem matches that used to obtain stable weights (Zubizarreta, 2015).
@@
$\sum^{n}_{i=1}\widehat{\alpha}_{i}\Phi((D_{i},Z_{i}))-\left(\sum^{n}_{i=1}\Big(\Phi((1,Z_{i}))-\Phi((0,Z_{i}))\Big)\right)=\bm{0}_{n},$
@@ 10.2 Logistic Models
$\alpha_{\beta}(X)=\mathbbm{1}[D_{i}=1]r_{\beta}((1,Z))-\mathbbm{1}[D_{i}=0]r_{\beta}((0,Z)),$
@@
$\pi_{\beta}(X)\coloneqq\frac{1}{1+\exp\big(-\beta^{\top}\Phi(Z)\big)}.$
and $\Phi\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a set of basis functions. Note that we do not include $D$ unlike the basis function in linear models. For this model, if we use the KL-divergence–flavored convex function defined in Section 6.3, we automatically attain covariate balancing, as discussed in Zhao (2019).

### v2

@@ 1.1 Contents of This Study
  • Generalized Riesz regression. We formulate the Riesz representer estimation problem as a Bregman divergence minimization problem. Bregman divergence depends on a convex function specified by us and yields various estimation methods by changing its definition. For example, Riesz regression and covariate balancing propensity scores are special cases. We refer to this estimation procedure as generalized Riesz regression or Bregman-Riesz regression.
@@ 6.2 Maximum Likelihood Estimation
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BR}}_{g^{\mathrm{KL}}}\big(\alpha\big)+\lambda J(\alpha),$
@@
$\widehat{\text{BR}}_{g^{\mathrm{KL}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\big|\alpha(X_{i})\big|-\log\big|\alpha((1,Z_{i}))\big|-\log\big|\alpha((0,Z_{i}))\big|\right).$
This corresponds to unnormalized KL (UKL) divergence minimization in density-ratio estimation (Sugiyama et al., 2011b), which generalizes Kullback-Leibler Importance Estimation. Procedure (KLIEP, Sugiyama et al., 2008; Nguyen et al., 2007).
@@ 6.3 Empirical Balancing
$g^{\mathrm{E}}(\alpha)=(|\alpha|-1)\log\left(\left|\alpha\right|-1\right)-|\alpha|.$
@@
$\partial g^{\mathrm{E}}(\alpha)=\begin{cases}-\log\left(-\alpha-1\right)&\text{if}\ \alpha<0\\ \log\left(\alpha-1\right)&\text{if}\ \alpha>1\end{cases}.$
@@
$\text{BR}_{g^{\mathrm{E}}}\big(\alpha\big)\coloneqq\mathbb{E}\Big[\log\big(|\alpha(X)|-1\big)+|\alpha(X)|-\log\big(\alpha((1,Z))-1\big)-\log\big(-\alpha((0,Z))-1\big)\Big].$
@@
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BR}}_{g^{\mathrm{E}}}\big(\alpha\big)+\lambda J(\alpha),$
@@
$\widehat{\text{BR}}_{g^{\mathrm{E}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\log\left(|\alpha(X_{i})|-1\right)+|\alpha(X_{i})|-\log\left(\alpha((1,Z_{i}))-1\right)-\log\left(-\alpha((0,Z_{i}))-1\right)\Big).$
@@
As in Section 6.2, model the Riesz representer via the inverse propensity score. Let $r((d,Z))$ model the inverse propensity score, that is, $r((1,Z))$ models $r_{0}((1,Z))\coloneqq 1/\pi_{0}(Z)$ and $r((0,Z))$ models $r_{0}((0,Z))\coloneqq 1/(1-\pi_{0}(Z))$. Based on this model, define
@@
Under this model, we write the Bregman divergence as
@@
Connection to the tailored loss.
Since $r((1,Z))=1/\pi(X)$ and $r((0,Z))=1/(1-\pi(X))$, we have $r((1,Z))-1=1/\big(r((0,Z))-1\big)$. Therefore, it holds that
@@ 7.2 Generalized Riesz Regression
With $g^{\text{KL}}(\alpha)=(|\alpha|-1)\log(|\alpha|-1)-|\alpha|$, the dual problem enforces moment balance between treated units and reweighted controls (automatic covariate balancing), aligning with entropy–balancing–style weights but learned through the Riesz objective.
@@ 9.2 Generalized Riesz Regression via Bregman Matching
which yields the classical LSIF used in Kanamori et al. (2009). While Chernozhukov et al. (2025) proposes Riesz regression for covariate shift adaptation, their Riesz regression problem is the same as LSIF, and their covariate shift adaptation method basically coincides with Kanamori et al. (2009) except for the use of regression functions. In parallel, Kato et al. (2024) also proposes a doubly robust form for covariate shift adaptation by extending Kanamori et al. (2009).
@@ 10.1 Linear Models
where $\Phi\colon\{1,0\}\times{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function. For this model, using the squared loss (Riesz regression) automatically attains covariate balancing, as discussed in Bruns-Smith et al. (2025).
@@
where $\bm{0}_{p}$ is the $p$-dimensional zero vector. This optimization problem matches that used to obtain stable weights (Zubizarreta, 2015).
@@
$\sum^{n}_{i=1}\widehat{\alpha}_{i}\Phi((D_{i},Z_{i}))-\left(\sum^{n}_{i=1}\Big(\Phi((1,Z_{i}))-\Phi((0,Z_{i}))\Big)\right)=\bm{0}_{p},$
@@ 10.2 Logistic Models
$\alpha_{\beta}(X)=\mathbbm{1}[D=1]r_{\beta}((1,Z))-\mathbbm{1}[D=0]r_{\beta}((0,Z)),$
@@
$\pi_{\beta}(X)\coloneqq\frac{1}{1+\exp\big(-\beta^{\top}\Phi(Z)\big)},$
and $\Phi\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function. Note that we do not include $D$ unlike the basis function in linear models. For this model, if we use the KL-divergence–flavored convex function defined in Section 6.3, we automatically attain covariate balancing, as discussed in Zhao (2019).